A 4-kg mass is at the origin and a 2-kg mass is at x=6 cm. Find the center of mass (x).
1
Expert's answer
2020-12-09T09:43:43-0500

"X_{CM}=\frac {m_1x_1+m_2x_2}{m_1+m_2}"


"X_{CM}=\frac {(4\times 0)+(2\times 6)}{2+6}"


"=\frac{12}{8}"


"=1.5 cm"